OpenStudy (anonymous):

Find the square root. of 25/49 please help and will be a fan!!

OpenStudy (durply):

5/7

OpenStudy (anonymous):

Thanks!!!

OpenStudy (durply):

Yup np

OpenStudy (durply):

anymore?

OpenStudy (anonymous):

yea let me check!!!

OpenStudy (anonymous):

Identify as rational or irrational. Explain. 13.001001001...

OpenStudy (durply):

A rational number is a number that can be expressed as a fraction. This means it can either be a terminating decimal or a recurring decimal. A terminating decimal is a decimal that doesn't recur (repeat for ever). A recurring decimal is a decimal that repeats a pattern of numbers after the decimal point. A good example of this is 1/3. 1/3 = 0.333333.... it is a rational number 13.001001001 is a rational number. Either you meant it to recur or not it makes no difference. 13.001001001 = 13 + 1001001/1000000000 as a terminating decimal 13.001001001....... = 12988/999 as a recurring decimal Examples of numbers that are not rational: sqrt(2), pi, exp they are irrational as they cannot be expressed as fractions.

OpenStudy (anonymous):

WOW thanks dude!!!

OpenStudy (durply):

ya np medal and fan plz

OpenStudy (anonymous):

So in shorts words the answer is ????

OpenStudy (durply):

XD

OpenStudy (anonymous):

What???

OpenStudy (durply):

read the last sentence

OpenStudy (anonymous):

kk thanks!!!

OpenStudy (durply):

medal and fan i can still help if needed

OpenStudy (anonymous):

Well a fan already so thanks a lot dude!!

OpenStudy (durply):

yup anything else?

OpenStudy (anonymous):

this last one please (Identify as rational or irrational. Explain. 4.7010010001...)

OpenStudy (durply):

n = 4.701001001001… 10000 n = 47010.01001001… : the 10000 takes us just past the 1st repeat 10 n = 47.01001001001… : the 10 takes us just to the 1st repeat subtract 9990 n = 46963 n = 46963 / 9990 rational

OpenStudy (anonymous):

Thanks Bro!!! laters!!!

OpenStudy (durply):

see ya what did you get?